OpenStudy (anonymous):

Solve 4x + 3 = 47 (1 point)

OpenStudy (midhun.madhu1987):

Subtract 3 from both sides.. what u get?

OpenStudy (anonymous):

Thank you so much!

OpenStudy (zbrandz23):

Subtract 3 on both sides, then divide by 4

OpenStudy (anonymous):

4x+3=47 4x=44 x=44/4 Final answer is x=11